sdftnstgs:

If BC = 6 and AD = 5, find DC.

imqwerty:

whats the whole question?

psirockin2:

let DC =x BD² = 5X according to pythagorean theorem: BD²+DC²=BC² (in triangle BCD) 5x+x²=6² x²+5x-36=0 x = 4 and x= -9 (according to quadratic formula) real solution is x=4 ( length cannot be negative) DC=4

mathmale:

"If BC = 6 and AD = 5, find DC" appears to be out of context. Where's the diagram that shows points B, C, A and D and line segments BC, AD and DC? Please share it.

celticcat:

yes. I'm baffled by the question as it stands!!

mathmale:

@sdftnstgs: Please repost your question when you're able to share the diagram. Thanks.

mathmale:

diagram, please? Use the Draw utility (below).